### v1

# Lagrangian Index Policy for

Abstract: We study the Lagrangian Index Policy (LIP) for restless multi-armed bandits with long-run average reward. In particular, we compare the performance of LIP with the performance of the Whittle Index Policy (WIP), both heuristic policies known to be asymptotically optimal under certain natural conditions. Even though in most cases their performances are very similar, in the cases when WIP shows bad performance, LIP continues to perform very well. We then propose reinforcement learning algorithms, both tabular and NN-based, to obtain online learning schemes for LIP in the model-free setting. The proposed reinforcement learning schemes for LIP requires significantly less memory than the analogous scheme for WIP. We calculate analytically the Lagrangian index for the restart model, which describes the optimal web crawling and the minimization of the weighted age of information. We also give a new proof of asymptotic optimality in case of homogeneous bandits as the number of arms goes to infinity, based on exchangeability and de Finetti’s theorem.
Key words: multi-armed bandits; restless bandits; Lagrangian index; Whittle index; exchangeability
1 Introduction
We consider restless multi-armed bandits with long-run average reward criterion as in [44]. These are Markov Decision Processes (MDPs) with two actions (‘active’ and ‘passive’) coupled only through the constraint on the number of active actions. This model has numerous applications in resource allocation [5, 26], queueing systems [1, 4, 18], web crawling and age of information [6, 23, 39], clinical trials and other health applications [12, 41, 42], just to name a few. See the books [22, 24, 37] and the recent survey [30] for thorough accounts of theory and applications of restless bandits.
The curse of dimensionality phenomenon makes the solution of the restless multi-armed bandit problem hard. In fact, the problem is known to be PSPACE-hard [33], which is a class among very hard problems in the computational complexity hierarchy. This motivated many researchers, starting from the seminal work of Whittle [44], to look for efficient heuristics. Whittle first relaxed the hard per stage constraint to an average constraint that requires only the asymptotic fraction of times the active mode was used to satisfy the constraint. This put the problem in the framework of classical constrained Markov decision processes, amenable to a Lagrange multiplier formulation. He used this to motivate the notion of subsidy, viz., a scalar $\lambda$ that is added to the running reward function when passive. He then defined indexability as the requirement that the set of states that are rendered passive under the optimal choice for the problem with subsidy, increase monotonically for each arm from the empty set to the entire state space as the subsidy is increased from $\infty$- ∞ to +$\infty$+ ∞ monotonically. If so, he assigned to each state an index, now called the Whittle index, equal to that value of subsidy for which both the active and passive actions are equally desirable for the given state. The heuristic policy he introduced, now called the Whittle Index Policy (WIP), is to order the observed states of the arms in decreasing values of the corresponding indices, settling ties according to some pre-specified rule, and then rendering top $M$ arms active, $M$ being the permitted number of active arms. While this prescription is arrived at via a relaxation of the original constraint, the resulting policy does satisfy the original per stage constraint. The priori nevertheless is a heuristic, but was shown to be asymptotically optimal by Weber and Weiss [43] (see also [40]) under certain conditions, more about these conditions later, when the number of arms goes to infinity. This and the fact that it works very well for many problems in practice has made it a very popular heuristic. The heuristic is also attractive because of its simplicity when the Whittle indices are explicitly known functions. There are many examples where they are, e.g., [6], [32], [39], [45]. See also the review article [30] for more examples.
That said, the problem still remains difficult when an explicit expression is unavailable, because the index needs a separate computation for each state, see, e.g., [7]. In the original Lagrangian formulation, there is only a single variable to be learnt, viz., the Lagrange multiplier. But exact solution of the Lagrangian formulation is optimal for the relaxed constraint, not the original per stage constraint. This suggests a heuristic akin to the Whittle index, but based on the actual Lagrange multiplier, viz., to rank states at each time according to the discrepancy between the ‘Q-values’ under active and passive actions under the Lagrange multiplier and render active the top $M$. This index was introduced in [16] in the finite horizon setting and mentioned in [20] as a possible heuristic for the infinite horizon with long-run average reward. We propose to call this new index the Lagrangian index. One more important advantage of the Lagrangian Index Policy (LIP) is that it does not require the indexability condition unlike the WIP.
Given that the Lagrangian index is computationally much more amenable when an explicit expression for the Whittle index is not available, this makes it an attractive alternative for applications. Reinforcement learning schemes for LIP that we present here requires significantly less memory than the corresponding scheme for WIP [7, 29, 31, 35, 36]. On the theoretical side, we also have a novel proof of asymptotic optimality for homogeneous arms as the number of arms goes to infinity, based on de Finetti’s theorem for exchangeable sequences of random variables. This may be of independent interest. Since LIP belongs to the family of LP-priority indices (see [40] for the definition and an extended treatment), the proof of asymptotic optimality of LIP follows from the results in [40] under the uniform global attractor assumption. Our proof also uses the uniform global attractor assumption. However, it can be easily carried out for both inequality and equality constraints.
The article is organized as follows. In the next section we formally introduce the model and define Whittle and Lagrangian indices. In Section 3 we present a tabular reinforcement learning algorithm for LIP and provide the proof of its convergence. Then, in Section 4 we describe a NN-based reinforcement learning algorithm for LIP and discuss its advantages with respect to the analogous algorithms for WIP. In Section 5 we calculate analytically the Lagrangian index for the restart problem, which can be used for the optimal web crawling and the minimization of the weighted age of information. In Section 6 we give an alternative proof of the asymptotic optimality of LIP. Finally, we provide numerical examples in Section 7 and conclude the paper with Section 8.
2 Model Formulation and Index Policies
Consider $N>1$ controlled discrete-time Markov chains (arms in bandit terminology) $X^{i}_{n},n\geq 0,$ for $1\leq i\leq N$, on finite state space $\mathcal{X}$, controlled by the $\mathcal{U}:=\{0,1\}$-valued control sequences $U^{i}_{n},n\geq 0$. Given the control $U^{i}_{n}$, the $i$-th chain moves from the current state $X^{i}_{n}\in\mathcal{X}$ to the next state $y\in\mathcal{X}$ according to the transition probability kernel $p^{i}(y|X^{i}_{n},U^{i}_{n})$ and obtains the reward $r^{i}(X^{i}_{n},U^{i}_{n})$. We seek to solve the following long-run average reward MDP problem:
($P_{0}$) Maximize the long-run average reward
$\lim_{n\to\infty}\frac{1}{n}E\left[\sum_{m=0}^{n-1}\sum_{i=1}^{N}r^{i}(X^{i}_{m},U^{i}_{m})\right],$ (2.1)
subject to the constraint
$\sum_{i=1}^{N}U^{i}_{n}=M,\quad\forall n\geq 0,$ (2.2)
for a prescribed $1\leq M<N$. The constraint (2.2) means that we need to set exactly $M$ arms active at each time step.
We shall consider the Whittle relaxation [44] of this problem, where we replace (2.2) by an average constraint. Thus the problem becomes
($P_{1}$) Maximize the average reward
$\lim_{n\to\infty}\frac{1}{n}E\left[\sum_{m=0}^{n-1}\sum_{i=1}^{N}r^{i}(X^{i}_{m},U^{i}_{m})\right],$ (2.3)
subject to the constraint
$\lim_{n\to\infty}\frac{1}{n}E\left[\sum_{m=0}^{n-1}\sum_{i=1}^{N}I\{U^{i}_{m}=1\}\right]=M,$ (2.4)
for a prescribed $1\leq M<N$.
[…]
($P_{2}$) Maximize the reward
$\lim_{n\to\infty}\frac{1}{n}E\left[\sum_{m=0}^{n-1}\sum_{i=1}^{N}\left(r^{i}(X^{i}_{m},U^{i}_{m})+\lambda_{*}\left(I\{U^{i}_{m}=1\}-\frac{M}{N}\right)\right)\right],$
where $\lambda_{*}$ is the Lagrange multiplier.
This separates the original MDP on $S^{N}$ into $N$ individual identical MDPs on $S$ that seek to maximize
$\lim_{n\to\infty}\frac{1}{n}E\left[\sum_{m=0}^{n-1}\left(r^{i}(X^{i}_{m},U^{i}_{m})+\lambda_{*}\left(I\{U^{i}_{m}=1\}-\frac{M}{N}\right)\right)\right],\quad i=1,...,N.$ (2.5)
This is possible because both the objective function and the constrained function are separable in the arms. We assume that these individual MDPs are irreducible under all stationary policies.
Define the ‘Q-values’ $Q^{i}(x,u),\ i=1,...,N,\ x\in S,\ u\in\mathcal{U}$, by
$Q^{i}(x,u)=ur^{i}(x,1)+(1-u)(\lambda_{*}+r^{i}(x,0))+\sum_{y}p^{i}(y|x,u)\max_{v}Q^{i}(y,v).$
Let us now define the Lagrangian index for the chains as follows:
$\gamma^{i}(x):=Q^{i}(x,1)-Q^{i}(x,0).$ (2.6)
This is the same as the ‘LP-index’ of [20], because the LP formulation and the Lagrangian formulation for $(P_{1})$ are equivalent. The Lagrangian index also belongs to the set of priority policies defined in [40].
Compare this with the Whittle index which is based on a subsidy $\lambda$ given for staying passive, and the index for state $x$ is the value of subsidy that makes the active and passive modes equally desirable. In particular, while its definition is the same as above, the Q-values in this case depend on the subsidy $\lambda$. If we denote this dependence explicitly by writing $Q^{i}_{\lambda}(x,u)$ for the Q-values, then the Whittle index for state $x$ is that value of $\lambda=\lambda(x)$ for which $\lambda$ is a solution of $Q^{i}_{\lambda}(x,1)-Q^{i}_{\lambda}(x,0)=0$. In contrast, the Lagrangian index for $x$ would correspond to $\gamma^{i}(x)=Q^{i}_{\lambda_{*}}(x,1)-Q^{i}_{\lambda_{*}}(x,0)$. This makes the computation of Lagrangian index much easier.
3 Tabular Q-Learning for Lagrangian Indices
In many practical situations the model ($p^{i}(y|x,u)$ and $r^{i}(x,u)$) is not known exactly or known approximately. This motivates us to develop model-free learning algorithms for the Lagrangian index. We propose two reinforcement learning algorithms: the first algorithm learns the Lagrangian index and does not enforce the hard constraint (2.2). This algorithm can be used for instance when a simulator is available. The second algorithm respects the hard constraint (2.2) during the learning process. Both algorithms are based on two-time scale primal-dual iterations for the saddle point reformulation of the Lagrangian relaxation [3, 10]:
$\lambda_{*}=\arg\min_{\lambda}\max_{\varphi}L(\varphi,\lambda),$ (3.1)
where
$L(\varphi,\lambda)=\lim_{n\to\infty}\frac{1}{n}E_{\varphi}\left[\sum_{m=0}^{n-1}\sum_{i=1}^{N}\left(r^{i}(X^{i}_{m},U^{i}_{m})+\lambda\left(I\{U^{i}_{m}=1\}-\frac{M}{N}\right)\right)\right]$
is the Lagrangian.
The primal iterates correspond to RVI Q-learning [2] and are carried out on the faster time scale. Namely, the Q-values are updated according to
$Q_{n+1}^{i}(x,u)=Q^{i}_{n}(x,u)+\alpha(\nu(x,u,n,i))I\{X_{n}=x,Z_{n}=u\}\times$
$\Big{(}ur(x,1)+(1-u)(\lambda_{n}+r(x,0))+\max_{v}Q^{i}_{n}(X_{n+1},v)-f(Q^{i}_{n})-Q^{i}_{n}(x,u)\Big{)},$ (3.2)
where $X^{i}_{n+1}\sim p^{i}(\cdot|x,u)$, $\nu(x,u,n,i)$ counts the number of times $Q^{i}(x,u)$ has been updated till time $n$, and where [^4] This is not the unique choice of $f(\cdot)$, see [2].
$f(Q)=\frac{1}{2|\mathcal{X}|}\sum_{i\in S}(Q(i,0)+Q(i,1)).$ (3.3)
The Lagrange multiplier is estimated by the dual iterates on the slower time scale:
$\lambda_{n+1}=\lambda_{n}-\beta(n)\left(\sum_{i=1}^{N}U^{i}_{n}-M\right),$ (3.4)
where $U^{i}_{n}$ is selected according to $\epsilon$-greedy policy, i.e.,
$U^{i}_{n}=\left\{\begin{array}[]{ll}\arg\max_{u}Q^{i}(X^{i}_{n},u),&\mbox{w.p.\ }1-\epsilon,\\ \mbox{Uni}(\mathcal{U}),&\mbox{w.p. \ }\epsilon.\end{array}\right.$ (3.5)
The time scale parameters need to satisfy the following conditions:
$\sum_{n}\alpha(n)=\infty,\quad\sum_{n}\alpha(n)^{2}<\infty,$
$\sum_{n}\beta(n)=\infty,\quad\sum_{n}\beta(n)^{2}<\infty,\quad\beta(n)=o(\alpha(n)).$ (3.6)
The first algorithm is summarized in Algorithm 1 and its convergence is established in the following theorem.
Theorem 3.1.
Under the conditions (3.6) for time steps, the iterates (3.2) and (3.4) converge a.s., i.e.,
$Q_{n}^{i}(x,u)\rightarrow Q_{\lambda_{*}}(x,u),$
and
$\lambda_{n}\rightarrow\lambda^{*}.$
as $n\rightarrow\infty$.
Proof.
(Sketch) The proof follows from standard ‘two time scale stochastic approximation’ argument (see, e.g., Chapter 8 of [15]). The $\{Q_{n}\}$ updates (3.2) operate on a faster time scale defined by the stepsizes $\{a(n)\}$ and the $\{\lambda_{n}\}$ iterates (3.4) operate on a slower time scale defined by the stepsizes $\{b(n)\}$. Thus the former see the latter as quasi-static, i.e., we can analyze (3.2) assuming $\lambda_{n}\approx$ a constant. Then (3.2) reduces to the RVI Q-learning algorithm of [2] and its convergence follows as in ibid. In turn, (3.4) sees the stationary expectation of $\sum_{i}U^{i}_{n}-M$. In fact, it is simply a stochastic subgradient descent scheme for the minimization of $\max_{\varphi}L(\varphi,\cdot)$. The map $L(\varphi,\lambda)$, as defined above, is affine in $\lambda$. Hence its maximum over $\varphi$ (albeit approximate here because of the $\epsilon$-greedy policy used) is convex in $\lambda$. Thus this is a stochastic subgradient descent for a convex function and converges a.s. to a global minimum. ∎
Algorithm 1 RL for Lagrangian index with relaxed constraint
1:initialize $\lambda$, Q(x,u) for all states and actions, $\epsilon$
2:for n = 1: $n_{end}$ do
3:     Choose action $U_{n}^{i}$ for each arm $i$ in $\epsilon$-greedy fashion according to (3.5)
4:     Update $X_{n+1}^{i}$ and reward $r_{n}^{i}$ from $X_{n}^{i}$ and $U_{n}^{i}$ for every arm $i$
5:     Update ($X_{n}^{i},U_{n}^{i}$)-th Q-value for each arm $i$ as in (3.2)
6:     Update common subsidy for passivity $\lambda$ as in (3.4)
7:     Optional: decrement $\epsilon$ as $\epsilon_{n+1}=max(0.01,\epsilon_{n}*0.99)$
8:end for
9:Calculate the new estimate of the index for each arm $i$:
10:      $\gamma^{i}_{n+1}(X^{i}_{n})=Q^{i}_{n}(X^{i}_{n},1)-Q^{i}_{n}(X^{i}_{n},0)$
In some cases a simulator may not be available and experiments with a real system require us to respect the hard constraint (2.2). In such cases, we propose a modification of Algorithm 1 described as Algorithm 2 by using virtual actions. The virtual actions are used to estimate the Lagrange multiplier and the actual actions are generated according to the Lagrangian indices. This way the algorithm respects the constraint on the resources even during the training phase.
Algorithm 2 RL for Lagrangian index with hard constraint
1:initialize $\lambda$, Q(x,u) for all states and actions, $\epsilon$
2:for n = 1: $n_{end}$ do
3:     Choose action $U_{n}^{i}$ according to $\epsilon$-Lagrangian policy, i.e.,
4:      w.p. $1-\epsilon$, activate top-$M$ arms according to Lagrangian index
5:      w.p. $\epsilon$, activate $M$ random arms
6:     Choose virtual action $\tilde{U}_{n}^{i}$ for each arm $i$ in $\epsilon$-greedy fashion as in (3.5)
7:     Update $X_{n+1}^{i}$ and reward $r_{n}^{i}$ from $X_{n}^{i}$ and $U_{n}^{i}$ for every arm $i$
8:     Update ($X_{n}^{i},U_{n}^{i}$)-th Q-value for each arm $i$ as in (3.2)
9:     Update common subsidy for passivity $\lambda$ using virtual actions, i.e.,
10:      $\lambda_{n+1}=\lambda_{n}-\beta(n)\left(\sum_{i=1}^{N}\tilde{U}^{i}_{n}-M\right)$
11:     Optional: decrement $\epsilon$ as $\epsilon_{n+1}=max(0.01,\epsilon_{n}*0.99)$
12:end for
13:Calculate the new estimate of the index for each arm $i$:
14:      $\gamma^{i}_{n+1}(X^{i}_{n})=Q^{i}_{n}(X^{i}_{n},1)-Q^{i}_{n}(X^{i}_{n},0)$
Theorem 3.2.
Under the conditions (3.6) for time steps, the iterates of Algorithm 2 converge a.s., i.e.,
$Q_{n}^{i}(x,u)\rightarrow Q_{\lambda_{*}}(x,u),$
and
$\lambda_{n}\rightarrow\lambda^{*}.$
as $n\rightarrow\infty$.
The proof is similar to that of Theorem 3.1.
We would like to note that Algorithms 1 and 2 are significantly simpler and use less memory than the reinforcement learning algorithm in [7] for the Whittle index. This is due to the fact that we do not need to use a reference state and many copies of Q-table.
4 Deep Q-Learning for Lagrangian Indices
Till now we have used tabular Q-learning for learning Q-values and ultimately the Lagrangian indices. Now, let us use the Deep Q-Learning Network (DQN) [27, 28] for approximation of the Q values. This will especially be useful in cases where the state space is large. A complete scheme of how our DQN algorithm works can be found in Algorithm 3. We note that in comparison with the DQN schemes for WIP [31, 35, 36], the DQN for LIP has a much simpler architecture and as a result numerically more stable and simpler to tune. This is because the DQN algortihms for WIP [31, 35, 36] require extra input for the reference state.
In the case of homogeneous arms, one-dimensional input to NN suffices, which represents the state of the arm. In the case of heterogeneous arms, a two-dimensional input to NN is used. The first dimension is used to give information about the type of arm, and the second dimension specifies the state of the arm. The neural network outputs are the Q-values for both possible actions. The actions are taken in an $\epsilon$-greedy fashion. Namely, the probability for exploration $\epsilon$ is initialized as one and then, at the end of every iteration, is reduced by the exploration-decay factor. The min value of $\epsilon$ is set to be 0.01.
The equation for the Q target used in training DQN is:
$Q_{target}(s_{n},a_{n})\leftarrow((1-a_{n})(r_{0}(s_{n})+\lambda_{n})+a_{n}r_{1}(s_{n})$
$+\underset{v\in\{0,1\}}{max}Q_{n}(s_{n+1},v)-\frac{1}{2d}\sum\limits_{k\in S}Q_{n}(k,0)+Q_{n}(k,1)).$ (4.1)
Note that in DQN we employ a second neural network, called the target model, for the computation of $\max_{v\in\{0,1\}}Q_{n}(s_{n+1},v)$. The reason for this decision is due to the maximisation bias of Q-learning: overestimating the Q-values results in the error increase over time, due to the target being $r+\gamma\max_{a}Q(s,a)$. The use of a second neural network for the target helps control this bias. This second neural network copies periodically the parameter values of the main network.
We store the $(s_{n}^{i},a_{n}^{i},r_{n}^{i},s_{n+1}^{i},\lambda_{n})$ tuples in an experience replay buffer, from which a batch of random samples is taken every time to train the main neural network. For each tuple $(s_{n}^{i},a_{n}^{i},r_{n}^{i},s_{n+1}^{i},\lambda_{n})$ we calculate Q target values using equation (4.1). Once we have calculated Q-values for all the examples in batch through $Q(\cdot)$ and its targets $Q_{target}(\cdot)$, we compute the loss function as the mean square error between the two and train the main neural network, through a standard Adam optimiser.
Algorithm 3 DQN Algorithm for Lagrangian Index
1:initialize $\lambda^{*}$, DQN policy, DQN target, Batch Size, $\epsilon$=1
2:for n = 1: $n_{end}$ do
3:     Update $\beta(n)$ as: $\beta(n)=\frac{1}{\lceil{nlog(n)\backslash 5000}\rceil+1}$
4:     for arm $i\in N$ do
5:         Choose action $U_{n}^{i}$ for each arm i in an $\epsilon$-greedy fashion
6:         Update $X_{n+1}^{i}$ and reward $r_{n}^{i}$ from $X_{n}^{i}$ and $U_{n}^{i}$ for every arm i
7:         Store $(X_{n}^{i},U{n}^{i},r_{n}^{i},X_{n+1}^{i},\lambda_{n})$ in experience replay buffer
8:         Train DQN policy model on one batch of $(X_{n}^{i},U_{n}^{i},r_{n}^{i},X_{n+1}^{i},\lambda_{n})$ tuples from experience replay buffer
9:     end for
10:     Update the DQN target model
11:     Update common subsidy for passivity for all arms $\lambda$ as: $\lambda_{n+1}=\lambda_{n}+\beta(n)(\sum\limits_{i=1}^{i=N}U_{n}^{i}-M)$
12:     Decrement $\epsilon$ as $\epsilon_{n+1}=max(0.01,\epsilon_{n}*0.99)$
13:end for
14:Calculate the new index for every state s of each arm i by:
15:      $\gamma^{i}(x)=Q_{\theta}(x,1)-Q_{\theta}(x,0)$
5 Application to Restart Model
Let us consider the following MDP, which can describe both the optimal web crawling [6] and the minimization of the weighted age of information [23, 38].
There are $N$ information sources, which can be viewed as arms of RMAB. The central device can probe $M$ information sources at a time. If the central device probes source $i$, it succeeds with probability $p_{i}$ (the failure can be due to either the distant server timeout or due to the loss of the probing packet). The freshness of the information at source $i$ at time slot $n$ is modelled by $X^{i}_{n}$ with the following dynamics:
$X^{i}_{n+1}=\left\{\begin{array}[]{ll}1,&\mbox{if \ }U^{i}_{n}=1\mbox{ \ and the probe is successful,}\\ X^{i}_{n}+1,&\mbox{otherwise}.\end{array}\right.$ (5.1)
For the relaxed problem formulation ($P_{2}$), we consider the following immediate reward
$r(x,u)=-w_{i}x+\lambda u,$ (5.2)
where $w_{i}$ gives the factor of importance of source $i$. We shall again use the saddle point reformulation (3.1) of the Lagrangian relaxation. Since we have moved the constraint to the Lagrangian, the optimal policy can be taken to be deterministic. In fact, it is proven in [23] (see also a similar development in [6]) that the optimal policy for the unconstrained MDP with the reward (5.2) is of threshold type [^5] The threshold property is usually proved by using additional structural properties such as monotonicity, convexity, submodularity, etc. of the value function. The needed properties are satisfied in the present model.. Therefore, we can search for an optimal policy in the form:
$\varphi^{i}_{*}(1|x)=\left\{\begin{array}[]{ll}0,&\mbox{ \ if \ }x<\bar{x}_{i},\\ 1,&\mbox{ \ if \ }x\geq\bar{x}_{i}.\end{array}\right.$ (5.3)
To analyse the model, we use the renewal argument dividing the time into cycles. A new cycle starts when $X^{i}_{n}=1$. The expected duration of cycle is given by
$E_{\varphi^{i}_{*}}[T^{i}_{cycle}]=\bar{x}_{i}-1+\frac{1}{p_{i}}.$ (5.4)
Let us calculate the average gain $g^{i}$ corresponding to the policy $\varphi^{i}_{*}$. By the renewal argument, we can write
$g^{i}=\frac{E_{\varphi^{i}_{*}}[R^{i}]}{E_{\varphi^{i}_{*}}[T^{i}_{cycle}]},$
where $E_{\varphi^{i}_{*}}[R^{i}]$ is the expected reward gained by arm $i$ on a cycle. (We suppress the dependence of $g$ on $\varphi^{i}_{*}$ for notational ease.) Next we calculate this by observing that a cycle can have two stages. Namely,
$E_{\varphi^{i}_{*}}[R^{i}]=E_{\varphi^{i}_{*}}[R^{i}_{1}]+E_{\varphi^{i}_{*}}[R^{i}_{2}],$
where
$E_{\varphi^{i}_{*}}[R^{i}_{1}]=-w_{i}\frac{(\bar{x}_{i}-1)\bar{x}_{i}}{2},$
corresponds to the first deterministic stage when $X^{i}_{n}<\bar{x}_{i}$, and
$E_{\varphi^{i}_{*}}[R^{i}_{2}]=E_{\varphi^{i}_{*}}\left[-w_{i}\frac{T_{2}(2\bar{x}_{i}+T_{2}-1)}{2}+\lambda T_{2}\right]$
$=-w_{i}\left(\frac{\bar{x}_{i}}{p_{i}}+\frac{1}{2}\frac{2-p_{i}}{p_{i}^{2}}-\frac{1}{2p_{i}}\right)+\frac{\lambda}{p_{i}}$
$=-w_{i}\left(\frac{\bar{x}_{i}-1}{p_{i}}+\frac{1}{p_{i}^{2}}\right)+\frac{\lambda}{p_{i}},$
where we used the fact that the duration of the second stage $T_{2}$ is geometrically distributed. Thus,
$g^{i}=\frac{-w_{i}\frac{(\bar{x}_{i}-1)\bar{x}_{i}}{2}-w_{i}\left(\frac{\bar{x}_{i}-1}{p_{i}}+\frac{1}{p_{i}^{2}}\right)+\frac{\lambda}{p_{i}}}{\bar{x}_{i}-1+\frac{1}{p_{i}}}.$
To optimize the above expression, we first find the minimizer of the continuous relaxation by equating to zero the derivative of $g^{i}$. This minimizer is given as the positive root of the following quadratic equation:
$p_{i}w_{i}\bar{x}_{i}^{2}+2w_{i}(1-p_{i})\bar{x}_{i}+(2\lambda-w_{i}(1-p_{i}))=0,$
which can be explicitly written as
$\tilde{x}_{i}=\frac{\sqrt{(1-p_{i})-2\lambda p_{i}/w_{i}}-(1-p_{i})}{p_{i}}.$
Since $\bar{x}_{i}$ is an integer, the maximal value of $g_{i}$ is given by
$g^{i}_{opt}=\max\{g^{i}(\lfloor\tilde{x}_{i}\rfloor),g^{i}(\lceil\tilde{x}_{i}\rceil)\},$ (5.5)
where $\lfloor a\rfloor$ is the integer part of $a$ and $\lceil a\rceil=\lfloor a\rfloor+1$. We shall denote by $\bar{x}^{i}_{opt}$ either $\lfloor\tilde{x}_{i}\rfloor$ or $\lceil\tilde{x}_{i}\rceil$ depending on whether the maximum above is attained at the former or at the latter.
Now, by (3.1), finding $\lambda$ is an easy one-dimensional convex optimization problem:
$\lambda_{*}=\arg\min_{\lambda}L(\phi_{opt},\lambda)=\arg\min_{\lambda}\sum_{i=1}^{N}g^{i}_{opt}(\lambda)-\lambda\alpha N.$ (5.6)
This problem can be efficiently solved for instance by the bisection method.
An interesting alternative to the solution of the above optimization problem is to estimate the Lagrange multiplier online as in (3.4). The advantage of this approach is that we effectively use the knowledge about the structure of the model and do not need to perform more demanding Q-updates (3.2).
Next, we need to satisfy the constraint (2.4). Again by the the renewal argument, we have
$\lim_{n\to\infty}\frac{1}{n}E_{\varphi^{i}_{*}}\left[\sum_{m=0}^{n-1}I\{U^{i}_{m}=1\}\right]=\frac{1}{E_{\varphi^{i}_{*}}[T^{i}_{cycle}]}=\frac{1}{\bar{x}_{i}-1+1/p_{i}}.$
In general, we shall have
$\sum_{i=1}^{N}\frac{1}{\bar{x}_{i,opt}(\lambda_{*})-1+1/p_{i}}\neq M.$
However, from the theory of the constrained MDP (see e.g., [34]), we know that there will be at least on arm $i_{*}$ such that
$g^{i_{*}}(\lfloor\tilde{x}_{i_{*}}\rfloor)=g^{i_{*}}(\lceil\tilde{x}_{i_{*}}\rceil).$
This allows us to do a randomization of the policy at that arm and to satisfy the constraint (2.4) exactly. We would like to note that in the case of a large number of arms, the discrepancy is very small and it makes a very small practical difference if we perform such randomization or just accept a minor violation of the constraint. Furthermore, as we see shortly, the expression of the Langrangian index depends only on the optimal gain and the optimal Lagrange multiplier and does not require the explicit expression of the optimal policy.
Specifying the expression (2.6), we can write
$\gamma^{i}(x)=Q^{i}(x,1)-Q^{i}(x,0)$
$=\left(-w_{i}x+\lambda_{*}-g^{i}_{opt}+p_{i}V^{i}(1)+(1-p_{i})V^{i}(x+1)\right)-\left(-w_{i}x-g^{i}_{opt}+V^{i}(x+1)\right)$
$=\lambda_{*}+p_{i}\left(V^{i}(1)-V^{i}(x+1)\right).$
Next we obtain $V$-values from the Bellman equation
$V^{i}(x)=\max_{u}\left\{-w_{i}x+\lambda_{*}u-g^{i}_{opt}+\sum_{y}p^{i}(y|x,u)V^{i}(y)\right\}.$
Since $V^{i}(\cdot)$ is the relative value, we can set $V^{i}(1)=0$ which simplifies calculations. We use the Bellman equation for two cases separately:
  1. 1.
    Case $x<\bar{x}_{i,opt}$: $V^{i}(x)=-w_{i}x-g^{i}_{opt}+V^{i}(x+1)$ or $V^{i}(x+1)=V^{i}(x)+w_{i}x+g^{i}_{opt}$. Consequently, we have
    $V^{i}(x)=V^{i}(1)+g^{i}_{opt}(x-1)+w_{i}\frac{(x-1)x}{2}=g^{i}_{opt}(x-1)+w_{i}\frac{(x-1)x}{2}.$
  2. 2.
    Case $x\geq\bar{x}_{i,opt}$: $V^{i}(x)=-w_{i}x+\lambda-g^{i}_{opt}+p_{i}V^{i}(1)+(1-p_{i})V^{i}(x+1)$, and hence $V^{i}(x+1)=(V^{i}(x)+w_{i}x-\lambda+g^{i}_{opt})/(1-p_{i})$. This is a linear difference equation, whose general solution is given by
    $V^{i}(x)=\frac{C_{i}}{(1-p_{i})^{x}}-\frac{w_{i}}{p_{i}}x+\frac{p_{i}(\lambda_{*}-g^{i}_{opt})-w_{i}(1-p_{i})}{p_{i}^{2}},$
    where the constant $C_{i}$ is found from the condition
    $V(\bar{x}_{i,opt})=g^{i}_{opt}(\bar{x}_{i,opt}-1)+w_{i}\frac{(\bar{x}_{i,opt}-1)\bar{x}_{i,opt}}{2}.$
Thus, we can summarize the above derivations in the following theorem and Algorithm 4 for online computation of the Lagrangian indices in the restart model. Note that in this version of the algorithm the probabilities $p_{i}$, $i=1,...,N$, are assumed to be known. However, a simple and efficient online estimator can be added if they need to be estimated as well. In fact, the problem of efficient estimation of change rates of information sources has been recently addressed in [8, 9]. We would like to emphasize that in the design of reinforcement learning algorithms it is always preferable to take into account the known structure of the model.
Theorem 5.1.
The Lagrangian index for the restart model has the following expression:
$\gamma^{i}(x)=\lambda_{*}-p_{i}V^{i}(x+1),$
where
$V^{i}(x)=g^{i}_{opt}(x-1)+w_{i}\frac{(x-1)x}{2},\quad x<\bar{x}_{i,opt},$
$V^{i}(x)=\frac{C_{i}}{(1-p_{i})^{x}}-\frac{w_{i}}{p_{i}}x+\frac{p_{i}(\lambda_{*}-g^{i}_{opt})-w_{i}(1-p_{i})}{p_{i}^{2}},\quad x\geq\bar{x}_{i,opt},$
and where the Lagrangian multiplier $\lambda_{*}$ is found from the one-dimensional convex optimization (5.6) and the optimal gain $g^{i}_{opt}$ is given by (5.5).
Algorithm 4 Algorithm for Restart Problem
1:initialize $\lambda=0$, $\epsilon=0.01$
2:for n = 1: $n_{end}$ do
3:     for arm $i\in N$ do
4:         Calculate $\tilde{x}_{i}=\frac{\sqrt{(1-p_{i})-2\lambda p_{i}/w_{i}}-(1-p_{i})}{p_{i}}.$
5:         Get $g^{i}_{opt}=\max\{g^{i}(\lfloor\tilde{x}_{i}\rfloor),g^{i}(\lceil\tilde{x}_{i}\rceil)\}$ using
6:$\;\;\;\;\;\;\;\;\;\;\;g^{i}=\frac{-w_{i}\frac{(x_{i}-1)x_{i}}{2}-w_{i}\left(\frac{x_{i}-1}{p_{i}}+\frac{1}{p_{i}^{2}}\right)+\frac{\lambda}{p_{i}}}{x_{i}-1+\frac{1}{p_{i}}}.$ substituting $x_{i}=\lfloor\tilde{x}_{i}\rfloor$ and
7:            $x_{i}=\lceil\tilde{x}_{i}\rceil$
8:         $\bar{x}^{i}_{opt}$ is either $\lfloor\tilde{x}_{i}\rfloor$ or $\lceil\tilde{x}_{i}\rceil$ depending on whether the maximum
9:           above is attained at the former or at the latter
10:         $Q^{i}(x,1)=-w_{i}x+\lambda-g^{i}_{opt}+(1-p_{i})V^{i}(x+1)$
11:$\;\;\;\;\;\;\;\;\;\;\;Q^{i}(x,0)=-w_{i}x-g^{i}_{opt}+V^{i}(x+1)$ where
12:$\;\;\;\;\;\;\;\;\;\;\;V^{i}(x)=g^{i}_{opt}(x-1)+w_{i}\frac{(x-1)x}{2},\quad x<\bar{x}_{i,opt},$
13:$\;\;\;\;\;\;\;\;\;\;\;V^{i}(x)=\frac{C_{i}}{(1-p_{i})^{x}}-\frac{w_{i}}{p_{i}}x+\frac{p_{i}(\lambda-g^{i}_{opt})-w_{i}(1-p_{i})}{p_{i}^{2}},\quad x\geq\bar{x}_{i,opt},$
14:         We get C from $\frac{C_{i}}{(1-p_{i})^{\bar{x}_{i,opt}}}-\frac{w_{i}}{p_{i}}{\bar{x}_{i,opt}}+\frac{p_{i}(\lambda-g^{i}_{opt})-w_{i}(1-p_{i})}{p_{i}^{2}}=$
15:           $V(\bar{x}_{i,opt})=g^{i}_{opt}(\bar{x}_{i,opt}-1)+w_{i}\frac{(\bar{x}_{i,opt}-1)\bar{x}_{i,opt}}{2}.$
16:         Choose action $a_{n}^{i}$ for each arm i in an $\epsilon$-greedy fashion based on
17:           $Q^{i}(x,1)$ and $Q^{i}(x,0)$
18:     end for
19:     Update $\lambda$ as: $\lambda_{n+1}=\lambda_{n}-\beta(n)(\sum\limits_{i=1}^{i=N}a_{n}^{i}-M)$
20:end for
21:$\gamma^{i}(x)=\lambda-p_{i}\left(V^{i}(x+1)\right)$
6 Asymptotic Optimality
In this section we provide an alternative proof of asymptotic optimality of the Lagrangian index for the case of homogeneous arms.
Since in this case the arms are identical, by symmetry one could divide both the displays in the statement of ($P_{1}$) by $N$ and reduce to a single constrained problem of maximizing the stationary expectation of $r(X_{m},U_{m})$ subject to the stationary expectation of $I\{U_{m}=1\}$ being equal to $\frac{M}{N}$. This symmetry implies in particular that $\lambda_{*}$ is also the Lagrange multiplier for this single arm constrained MDP. This MDP, however, is a complete artifice, but it will play a role in our proof. We assume that these MDPs are irreducible under all stationary policies. From the theory of constrained MDPs [3, 11, 34], we have the following.
Theorem 6.1.
For a single constraint MDP, there exists an optimal stationary randomized policy $\varphi_{*}(\cdot|\cdot)$ that needs to randomize in at most one state, between at most two control choices.
Since $|\mathcal{U}|=2$ in the present case, the latter ‘at most’ means that it needs to randomize at most in one state, say $x_{*}$, between 00 and $1$ with probabilities $1-p,p$ resp., for some $p\in[0,1]$. Let $\mathcal{A}:=\{x\in S:\varphi_{*}(1|x)=1\}$ denote the set of states that are always active under this policy. Let $\pi_{*}$ denote the unique stationary distribution of $\{X_{n}\}$ under the optimal policy. Since the constraint (2.4) is met under the optimal policy $\varphi_{*}$, we have
$\pi_{*}(\mathcal{A})+p\pi_{*}(\{x^{i}_{*}\})=\alpha:=\frac{M}{N}.$ (6.1)
In the following, we shall compare two policies:
  1. 1.
    the relaxed optimal policy, denoted by OPT and given by $\varphi^{*}$, and,
  2. 2.
    the Lagrange index policy LI that renders active at time $n$ the chains $i,1\leq i\leq N,$ with the top $M$ values of $\gamma(X^{i}_{n})$, breaking any ties with equal probability.
We are interested in the $N\uparrow\infty$ limit.
Recall the definition of $\alpha$ in the original set up. We take $\alpha$ as a prescribed fraction in $(0,1)$ and define:
$M_{N}:=$ the smallest integer exceeding $\alpha N$,
$m_{N}:=M_{N}-1=$ the largest integer not exceeding $\alpha N$.
Replace $M$ by $M_{N}$, resp., $m_{N}$, in the foregoing. Clearly,
$\alpha_{N}:=\frac{M_{N}}{N}\to\alpha,\ \alpha_{N}^{\prime}:=\frac{m_{N}}{N}\to\alpha\ \mbox{as}\ N\uparrow\infty.$
Under OPT, the chains are independent of each other. Let $\epsilon>0$. Under OPT, (6.1) and the strong law of large numbers imply that
[…]
for all $n\geq 0$.
Under LI, the chains are not independent, but they form an exchangeable family. This is because, by our choice of the index rule and the tie-breaking protocol, their joint laws are invariant under permutations. We shall work directly with $N=\infty$.
Let $S$ denote the state space. Number the states according to decreasing values of the Lagrange index $\gamma(j):=Q_{\lambda^{*}}(j,1)-Q_{\lambda^{*}}(j,0)$. Let
$S^{+}$ ::=:= $\{j\in S:\gamma(j)>0\},$
$S^{0}$ ::=:= $\{j\in S:\gamma(j)=0\},$
$S^{-}$ ::=:= $\{j\in S:\gamma(j)<0\}.$
Then $|S^{0}|\leq 1$ by Theorem 6.1. The following is immediate from Theorem 6.1.
Theorem 6.2.
For a single chain, the optimal control will be assigned as per the following procedure:
Procedure 1: Assign $1$ for $j\in S^{+}$, 00 for $j\in S^{-}$, and $1$ with probability $\beta$ and 00 with probability $1-\beta$ for $i\in S^{0}$. Here $\beta\in[0,1]$ is such that if $\pi$ denotes ths stationary distribution of the chain, then $\pi(S^{+})+\beta\pi(S^{0})=\alpha$.
Let $X^{i}_{n},1\leq i<\infty,$ denote the current states of the arms. List the arms as $1,2,...,$ according to decreasing values of $\gamma(X^{i}_{n})$’s. We follow the following procedure, which is the counterpart of ‘Procedure 1’ above for the infinite sequence:
Procedure 2: Club together the $X^{i}_{n}$’s with equal indices into a single group. Rank these groups according to decreasing $\Lambda^{*}(X^{i}_{n})$. Assign control $1$ to the top $\ell^{*}-1$ such groups, where
$\ell^{*}:=\min\{k\geq 1:$ the total fraction of chains with states in group $k$, $k\leq\ell^{*}$ is $\geq\alpha\}$.
For chains in state $\ell^{*}$, a fraction $\kappa:=(\alpha\ -$ the total fraction of chains in states in group $j$, $j<\ell^{*})$ is assigned control $1$, the rest 00, in an exchangeable manner.
The above procedure is in fact deterministic, because the fraction of arms in each state is deterministic by Lemma 1, p. 6, of [20]. In fact, let $\nu_{n}\in\mathcal{P}(S)$ denote the deterministic distribution of arms across the state at time $n$, i.e., the fraction of arms in state $i$ is $\nu_{n}(i),i\in S$. Then written as a row vector, $\{\nu_{n}\}$ evolve as $\nu_{n+1}=\nu_{n}P,n\geq 0,$ where $P$ is the common transition matrix of the chains. Then the ‘Global Attractor Hypothesis’ from [40], as enunciated in [20], is:
()($\dagger$)( † ) For any choice of $\nu_{0}\in\mathcal{P}(S)$, $\nu_{n}\to$ a unique $\nu^{*}\in\mathcal{P}(S)$ satisfying $\nu^{*}=\nu^{*}P$.
In particular, ()($\dagger$)( † ) implies that the limiting deterministic distribution $\nu^{*}$ across states is uniquely specified across time and therefore the above procedure is well-defined and deterministic.
Theorem 6.3.
The arms are asymptotically independent.
Proof.
Let $\mathcal{E}$ denote the $\sigma$-field of exchangeable events. By de Finetti’s theorem [25], conditioned on $\mathcal{E}$, $X^{i}_{n}$’s are independent chains. Thus conditioned on $\mathcal{E}$, the controls are assigned to each arm independently, a.s. Let $A$ denote the event that Procedure 2 is applied sample point wise. Since $P(A)=1$, we also have $P(A|\mathcal{E})=1$ a.s., i.e., the above procedure holds a.s. for the independent chains. Under the global attractor hypothesis, the asymptotic fraction of arms in any given state is deterministic, collectively described by the probability vector $\nu^{*}$. Let $\nu^{\dagger}(\omega)\in\mathcal{P}(S)$ denote the asymptotic fraction of arms in different states for the conditional chain conditioned on $\mathcal{E}$, where $\omega$ denotes the sample point. In particular, $\omega\mapsto\nu^{\dagger}(\omega)$ is $\mathcal{E}$-measurable. Then $\nu^{*}$ is perforce a mixture thereof, i.e., $\nu^{*}=E[\nu^{\dagger}(\omega)]$.
Note that Procedure 2 above is deterministic except for the randomization for the state $\ell^{*}$. Since $P(A|\mathcal{E})=1$ a.s., this property must hold for the aforementioned collection of independent chains, a.s.-$\omega$. Furthermore, since the chains are independent, the randomization must be independently done. The fact that $A$ holds conditionally (a.s.) also implies that the asymptotic fraction of active arms must remain $\alpha$, a.s. $\omega$. It is easy to see then that the only way all these conditions can be satisfied is if Procedure 1, which is optimal for a single chain, is followed for each chain (arm), whence asymptotic optimality follows under the global attractor hypothesis. ∎
Remark 6.4.
De Finetti’s theorem does not hold for finite exchangeable sequences. An approximate version of the theorem is available [17] and could be made a basis of an analogous argument that begins with $N<\infty$ arms and then passes to the $N\to\infty$ limit. However, that extra effort seems unwarranted.
7 Numerical Examples
7.1 Restart Model
In this first numerical experiment, we consider the restart model as described in Section 5 and apply Algorithm 4. There are four different categories of arms:
  • Type 1: $p_{1}=0.95,w_{1}=0.9$ (reliable, important)
  • Type 2: $p_{2}=0.95,w_{2}=0.2$ (reliable, not important)
  • Type 3: $p_{3}=0.7,w_{3}=0.95$ (not very reliable, important)
  • Type 4: $p_{4}=0.7,w_{4}=0.2$ (not very reliable, not important)
A total of 100 arms are taken, with 25 belonging to each category and the budget is set as 16 arms. Therefore, in this example $N=100$ and $M=16$. Note that for all the 100 arms we still learn just one common Lagrangian multiplier.
Figure 1(a) shows the convergence of the subsidy (Lagrange multiplier) to the value -11.6. In the right Figure 1(b) we plot the value of the optimal Lagrangian function (optimized with respect to the policy) at different values of the Lagrange multiplier and as expected we see that the Lagrangian function is piecewise linear and takes the minimum value around -11.6 (shown by the red line) same as the value of the Lagrange multiplier we obtain by our Algorithm 4.
(a) Subsidy (Lagrange multiplier) estimation.
(b) Optimal Lagrangian.
Figure 1: Restart Model. Subsidy (Lagrange multiplier).
In Figure 2 we compare the performances of LIP with WIP. We used the Whittle indices calculated in [38]. As we see in the restart example LIP performs as good as WIP.
Figure 2: Restart Model. LIP and WIP average reward comparison.
7.2 A Non Whittle-Indexable Problem
In this example, we test the performance of the Lagrangian Index in a Non-Whittle Indexable Problem as described in Appendix A.4 of [19].
The transition probability matrices for this problem are given by
$P_{0}=\begin{bmatrix}0.005&0.793&0.202\\ 0.027&0.558&0.415\\ 0.736&0.249&0.015\end{bmatrix}\quad\text{and}\quad P_{1}=\begin{bmatrix}0.718&0.254&0.028\\ 0.347&0.097&0.556\\ 0.015&0.956&0.029\end{bmatrix}$
and the rewards are given by
$r(1,0)=0,\quad r(2,0)=0,\quad r(3,0)=0,$
$r(1,1)=0.699,\quad r(2,1)=0.362,\quad r(3,1)=0.715.$
In many practical cases even when the problem is Non Whittle indexable, Whittle indices are used empirically. Hence, we compute online the Whittle indices for this problem by the algorithm from [7] and compare the performance of Lagrangian indices against them. We test both Algorithm 1 and Algorithm 2 on this example, see Figure 3 and Figure 4, respectively. As seen from the plots, the Lagrangian indices result in a much higher reward as compared to the reward obtained by the algorithm estimating Whittle indices. Also, we note that even though Algorithm 2 respects the hard constraint and uses virtual actions to estimate subsidy, this does not affect the rate of convergence of the moving average reward.
(a) Subsidy (Lagrange multiplier) estimation.
(b) Average reward.
Figure 3: Non Whittle Indexable problem. Algorithm 1.
(a) Subsidy (Lagrange multiplier) estimation.
(b) Average reward.
Figure 4: Non Whittle Indexable problem. Algorithm 2.
7.3 Deadline Scheduling Problem
Let us now consider the deadline scheduling problem studied in [45]. The states of this problem are formed by two different variables: the service time $B\in[0,9]$ and the deadline $T\in[0,12]$, leading to a total of $|S|=130$ states (although several of these states are not accessible by the underlying Markov chain). The variable $B$ represents the amount of workload pending to complete a certain job while the variable $T$ represents the remaining time available to perform it. When there is no job at the $i^{th}$ position, the state is $(0,0)$, while otherwise it is $(T_{n}^{i},B_{n}^{i})$. In each iteration, the state transition from $s_{n}^{i}=(T_{n}^{i},B_{n}^{i})$ depends on the action $a_{n}^{i}$ in the following way:
$s_{n+1}^{i}=\begin{cases}(T_{n}^{i}-1,(B_{n}^{i}-a_{n}^{i})^{+}),&\text{if }T_{n}^{i}>1,\\ (T,B)\text{ with prob. }q(T,B),&\text{if }T_{n}^{i}\leq 1.\end{cases}$ (7.1)
where $b^{+}=\max(b,0)$. When $T$ = 1, the deadline for completing the job ends, and the new state (including the empty state $(0,0)$) is chosen uniformly at random. The value of $B$, the workload to be completed, is only reduced if the action is active. If the scheduler reaches the state ($T=1$, $B>0$), the job cannot be finished on time, and an extra penalty $F(B_{n}^{i}-a_{n}^{i})=0.2(B_{n}^{i}-a_{n}^{i})^{2}$ is incurred. In addition, activating the arms involves a fixed cost $c=0.8$. The rewards are given by
$r_{n}^{i}(s_{n}^{i},a_{n}^{i},c)=\begin{cases}(1-c)a_{n}^{i},&\text{if }B_{n}^{i}>0\text{ and }T_{n}^{i}>1,\\ (1-c)a_{n}^{i}-F(B_{n}^{i}-a_{n}^{i}),&\text{if }B_{n}^{i}>0\text{ and }T_{n}^{i}=1,\\ 0,&\text{otherwise.}\end{cases}$ (7.2)
The authors of [45] provide an explicit expression for Whittle index, given by:
$\lambda(T,B,c)=\begin{cases}0,&\text{if }B=0,\\ 1-c,&\text{if }1\leq B\leq T-1,\\ F(B-T-1)-F(B-T)+1-c,&\text{if }T\leq B.\end{cases}$ (7.3)
As the state space is large we use DQN to learn the Lagrangian indices. A neural network with 3 hidden layers having (512, 256, 128) neurons connected through the ReLU activation is used. A batch size of 32 is used and the size of the experience replay buffer is 1000. The exploration decay factor is 0.9995 and the learning rate for the Adam optimiser is $10^{-5}$.
We first consider the homogeneous case using $N$ = 5 and $M$ = 2, with a processing cost $c=0.8$. The results are presented in Figure 5.
(a) Subsidy (Lagrange multiplier) estimation.
(b) Average reward.
Figure 5: Deadline scheduling problem (homogeneous arms). Algorithm 3.
Then we consider the scenario of heterogeneous arms, where each arm has a state space of size $|S|=130$ and $N=20$. The results are presented in Figure 6. To differentiate the arms, we assign different activation cost values $c=\{0.1,0.3,0.6,0.8\}$ for every 5 arms. As was suggested in Section 4, we take into account heterogeneity by increasing the dimension of the input to the neural network. The first dimension is used to give information about the type of arm and the other dimensions represent the state of the arm. Specifically, the values of the first input are $0,1,2,3$ for $c=0.1$, $c=0.3$, $c=0.6$ and $c=0.8$.
(a) Subsidy (Lagrange multiplier) estimation.
(b) Average reward.
Figure 6: Deadline scheduling problem (heterogeneous arms). Algorithm 3.
Here, as was the case in the restart model, the performance of LIP is comparable to the performance of WIP. However, since we do not need to use a reference state in our algorithms for LIP as opposed to the algorithms from [7, 31, 35, 36], the implementation of the reinforcement learning algorithms for LIP is much more straightforward than that for WIP and at the end uses less number of computer operations. This leads to the gain in time and memory.
8 Conclusion
We have proposed and analyzed the Lagrangian index for average cost restless bandits. It is computationally more tractable than Whittle index when the latter is not explicitly known and does not require Whittle indexability. We have calculated analytically the Lagrangian index for the restart model. Then, we have provided several numerical experiments. LIP is seen to perform comparably well on problems that are Whittle indexable and as well or better on problems that are not. We also prove its asymptotic optimality in the infinite arms limit under the popular ‘global attractor hypothesis’ using a novel argument based on exchangeability and de Finetti’s theorem.
Acknowledgements
The authors would like to thank E. Altman and A. Piunovskiy for helpful discussions about constrained Markov decision processes.
Statements and Declarations
The authors declare no competing or conflicting interests related to the present work.